OpenStudy (anonymous):

Solve the system y = 2x + 6 and 3y = 6x + 18 by using graph paper or graphing technology. What is the solution to the system? No Solutions (3, 6) Infinite Solutions (2, 6) Solve the system y = -x + 6 and y = x - 8 by using graph paper or graphing technology. (1.5 , -6.5) (7 , -1) (10 , 2) (2 , 2) Help me understand these two? Thanks

OpenStudy (anonymous):

There are infinite solutions to the first question since they are the same line. \[3y = 6x+18 \rightarrow y = 2x+6\]

OpenStudy (anonymous):

For the second , just equate the two and get the value of x \[-x+6 = x-8\]

OpenStudy (anonymous):

Wait a second, you want me to do that equation? That'd give me x?

OpenStudy (anonymous):

yes once u get x, then u can plug the value of x in either of the two equation to get y

OpenStudy (anonymous):

...Wheres the pen and paper

OpenStudy (anonymous):

answer is 7,1-